# Wave pulses with unusual asymptotical behavior at infinity

### v2

@@ Abstract
The behavior of wave signals in the far zone is not only of theoretical interest but also of paramount practical importance in communications and other fields of applications of optical, electromagnetic or acoustic waves. Long time ago T. T. Wu introduced models of “electromagnetic missiles” whose decay could be made arbitrarily slower than the usual inverse distance by an appropriate choice of the high frequency portion of the source spectrum. Very recent work by Plachenov and Kiselev introduced a finite-energy scalar wave solution, different from Wu’s, decaying slower than inversely proportional with the distance. A physical explanation for the unusual asymptotic behavior of the latter will be given in this article. Furthermore, two additional examples of scalar wave pulses characterized by abnormal slow decay in the far zone will be given and their asymptotic behavior will be discussed. A proof of feasibility of acoustic and electromagnetic fields with the abnormal asymptotics will be described.
wave equation, unidirectional wavepackets, ultrashort electromagnetic and acoustical pulses
preprint: APS/123-QED
I Introduction
It is a textbook truth that solutions of the wave equation decay inversely proportional to distance in the far field or the wave zone (see, e.g., [1, 2]). However, a long time ago Wu [3] showed that the decay can be made arbitrarily slower by a careful choice of the high-frequency wing of the source spectrum. In the 1980s, closely related to Wu’s ”electromagnetic missiles”, another subject dealing with uncommon solutions of the wave equation emerged—the so-called localized or non-diffracting or space-time wave packets possessing exotic properties, such as self-healing, superluminality, and invariant propagation without spread or decay, theoretically up to infinity. By now this subject has grown into a large research field of its own, see monograph [4] and reviews [5, 6, 7, 8, 9]. Therefore, the decay in the wave zone as a specific subject in mathematical physics has remained somewhat in the shadows. However, in recent years, several papers on the study of asymptotical behavior have appeared, see [10, 11, 12] and Refs. therein.
The very recent paper by Plachenov and Kiselev [12] introduced a finite-energy solution, different from Wu’s model [3], that decays slower than inversely proportional to the distance.
They consider the following simple solution of the homogeneous wave equation (we have somewhat changed the designations)
@@ I Introduction
where $R=\left(x^{2}+y^{2}+z^{2}\right)^{1/2}$ is the distance of the field point from the origin, $c$ is the velocity of light (or sound), and $t_{s}$ is a constant that determines the width of the pulse. $\psi\left(R,t\right)$ is a special case of a splash pulse [13]. It is frequently used for modeling ultrashort pulses and constitutes a spherically propagating pulse, while the spherical surface, where the pulse peaks at $R\approx ct,$ collapses into the origin at negative times $t<0$ and expands for $t>0.$ The real part $\operatorname{Re}\psi\left(R,t\right)$ represents a bipolar pulse and the imaginary part $\operatorname{Im}\psi\left(R,t\right)$ a unipolar one everywhere where $R\gg ct_{s}$, and vice versa near the origin where $R<ct_{s}$.
@@
The limit of the type of Eq. (2) plays an important role in various fields of mathematical physics and it is known that for finite-energy ”well-behaving” solutions of the wave equation the limit $L$ should be finite (incl. $L=0$), see, e.g., [1, 2, 10, 11].
@@
[…]
in contradistinction to the usual behaviour of fields in the wave zone, the quantity in Eq. (2) is diverging:
@@
They also showed that $\Psi\left(R,t\right)$ has finite total energy, i.e., it is not an unphysical solution.
The purpose of our article is to present additional examples of pulses with unusual asymptotics and to reveal physical causes of abnormally slow decay. In the next section we consider one of such examples—the primitive (antiderivative) of the wave function of a unidirectional pulse. Sect. III is devoted to another example—the fractional splash pulse which already itself, without taking its time integral, decays very slowly in the far zone.
The authors in the short communication [12] neither illustrate graphically the behavior of $\Psi\left(R,t\right)$ nor interpret physically its unusual asymptotics. In Sec. IV we try to fill in this gap. In Sect. V we discuss some subtleties of the slow decay of the other pulses studied in Sect. II and III and consider the physical feasibility in optics, electromagnetics, or acoustics of the pulses considered in this study.
II Hemispherical unidirectional pulse
@@
Similarly to $\psi\left(R,t\right)$ in Eq. (1), the real and imaginary parts of $u\left(R,t\right)$ represent bipolar and unipolar pulses, but in contrast to the spherically symmetric case, in the given case the pulses ”ride” on a collapsing-expanding tube-like surface with a hemispherical surface inside (see plots below).
@@
But how does its primitive behave? Taking the integral of $u\left(\rho,z,t\right)$ is a complicated task. It can be found neither in tables of integrals nor can be taken by the newest versions of computation packages Mathematica and Maple. The integral over time from $\infty$- ∞ to $t$ was taken first by setting $t_{s}=z_{s}=0$ and then making replacements $z\rightarrow z-iz_{s}$ and $t\rightarrow t+it_{s}$. After some algebra, we found the following expression for the primitive, where $g\left(\rho,t\right)$ is given in Eq. (5).
@@
$U_{1}\left(\rho,z,t\right)$ $\equiv\ln\frac{ct+ict_{s}+h\left(\rho,z\right)}{ct+ict_{s}-h\left(\rho,z\right)},\qquad h\left(\rho,z\right)\equiv\sqrt{\rho^{2}+\left(z-iz_{s}\right)^{2}},$
$U_{2}\left(\rho,z,t\right)$ $\equiv\ln\frac{\left[\left(-ict+ct_{s}\right)\left(z-iz_{s}\right)+g\left(\rho,t\right)h\left(\rho,z\right)\right]\left[z-iz_{s}-h\left(\rho,z\right)\right]}{\left[\left(-ict+ct_{s}\right)\left(z-iz_{s}\right)-g\left(\rho,t\right)h\left(\rho,z\right)\right]\left[z-iz_{s}+h\left(\rho,z\right)\right]}.$
The behavior of $U\left(\rho,z,t\right)$ is depicted in Figs. 1 and 2. We see that the pulse peaks on the surface of a collapsing-expanding tube with a hemisphere inside it, while the radius of both is equal to $|ct|$. Hence, $U\left(\rho,z,t\right)$ behaves similarly to $u\left(\rho,z,t\right)$ from Eq. (5), the most remarkable difference being that the peak of the real part of the latter is abruptly and symmetrically bipolar, except at $t=0$ when it is unipolar contrary to Fig. 1.
@@
Another exception where the limit proves to be finite and equals to $\ln 2$ is when time $t$ runs to $\infty$- ∞ but $z$ to +$\infty$+ ∞. In both these exceptional cases the pulse peak is not involved into the process, as can be seen by looking at Figs. 1 and 2.
We calculated the Poynting vector and energy density of $U\left(\rho,z,t\right)$ according to known formulas (see, e.g., Eqs. (1)-(5) of Ref. [18]) for scalar fields, as well as the total energy. The latter is finite. The former two have normal asymptotics in the sense of finite values of the limit of the type in Eq. (2) as they have to be since both consist of derivatives of $U\left(\rho,z,t\right)$ which have normal asymptotics.
Despite the unidirectionality of $U\left(\rho,z,t\right)$—integration as a linear operation preserves the unidirectional spectrum of $u\left(\rho,z,t\right)$— in certain spatial regions the energy flux turned out to be directed opposite to the azis $z$, and at some points the velocity of such an energy backflow even reached almost the value $-c$.
@@
Finally, we found that the imaginary part of $U\left(\rho,z,t\right)$ is a ”strange” field [20, 21, 22] in the sense that in certain regions the time integral from $\infty$- ∞ to +$\infty$+ ∞ is not zero. This is understandable, because the imaginary part is a unipolar pulse.
III Fractional splash pulse
Members of a general family of splash pulses are derived as spectral superpositions of the most investigated localized propagation-invariant unipolar pulse—the focus wave mode [5, 6, 4]. A specific fractional splash pulse is given by the following expression
@@ III Fractional splash pulse
×$\times\frac{1}{\left(a_{2}-i\left(z+ct\right)+\rho^{2}\left(a_{1}+i\left(z-ct\right)\right)^{-1}\right)^{3/4}},$
where $a_{1}$ and $a_{2}$ are positive parameters. The general form of the splash mode pulse is given by the expression in Eq. (13), with the power $3/4$ in the denominator replaced by $\nu+1$, with $\nu>-1$. The fractional pulse above corresponds to $\nu=-1/4$.
The behavior of $f\left(\rho,z,t\right)$ is depicted in Fig. 3. Both the real and imaginary parts of the pulse possess a collapsing-expanding spherical structure, but a peak appears only in a narrow cone of one hemisphere (of positive $z$ if $ct>0$ and of negative $z$ if $ct<0$). The real part has mirror symmetry with respect to time reversal $t\rightarrow-t$, while the imaginary part additionally reverses its sign. Outside the region of the origin, the pulse tops of the real part and the imaginary part do not coincide: the top of the former is shifted farther from the origin $z=0$ than the point $z=ct$ by a generally small distance about $32a_{1}/$ 3, and the top of the imaginary part is shifted closer by the distance about $a_{1}/4$. This follows from a detailed inspection of Eq. (13), as well as from plots which will be presented in Section V. But the most remarkable distinction from the pulses considered in the previous sections is that the fractional splash pulse itself, without the need to take its antiderivative, exhibits abnormal asymptotics as the limits of the type in Eq. (2) show. The limits along different directions turn out to be as follows.
@@
Eq. (14) shows that the peaks of both the real and imaginary parts decay slower than the normal decay $\sim 1/z$ when one moves with the pulse top along the positive direction of the axis $z$ towards infinity. How they approach infinity will be discussed in Sect. V. As was mentioned above, the tops of the pulses of the real and imaginary parts of $f\left(\rho,z,t\right)$ are shifted from the point $z=ct$, and therefore the value of $\Delta$ affects differently the limits of the parts. If $ct>0$ and $z>0$, a simple choice is $\Delta=0$ in which case the running point is simultaneously close to the tops of the pulses of both the real and imaginary parts and the exponent in Eq. (14) shows that at this point both $ct\,\Re f\left(\rho,z=ct,t\right)$ and $ct\,\Im f\left(\rho,z=ct,t\right)$ approach infinity in the same way.
@@
We calculated the Poynting vector and energy density of $f\left(\rho,z,t\right)$. Although individually time and space derivatives of the fractional splash pulse exhibit abnormal asymptotic behavior, the energy density and the Poynting vector when inserted into Eq. (14) instead of $f$ do not result in an infinite limit because both quantities are quadratic in time and space derivatives. However, for energy-type quantities, a normal decay is inversely proportional to the square of the distance. This means that in testing the asymptotic behavior, $ct$ must be replaced by $c^{2}t^{2}$ in the expression of the limit. In doing so, we arrive at the conclusion that the energy density and the Poynting vector also decay abnormally slowly. The total energy is finite. $f\left(\rho,z,t\right)$ is square integrable, which is also important from a physical point of view as will be discussed in Sect. V.
We studied also other fractional splash pulses which are given by the same expression in Eq. (13) but with other values of the power instead of $3/4$ in the denominator. If the power is $1/2$, the pulse has the same properties as $f\left(\rho,z,t\right)$, but unfortunately its total energy is infinite. But if the value of the power is between $1/2$ and $1$, i.e., if $\frac{1}{2}<\nu+1<1$, the pulse has slower than normal decay and at the same time its wave function is square-integrable and has finite total energy.
IV Causes of uncommon asymptotical behavior at infinity
@@
$\rho(\textbf{r},t)=\delta(\textbf{r})\hskip 2.0pt\frac{1}{2c(t+it_{s})}~{},$ (19)
@@
In Fig. 4 the imaginary parts of the expanding and converging waves given by Eq. (21) and Eq. (22), respectively, are plotted against the radius $R$ at three different instants of time. The curves are plotted with reverse sign for better readability and are scaled—multiplied by $R$—in order to eliminate the factor $1/R$. Thanks to the scaling we observe that the peaks of the pulses at different time instants are of equal height. This correponds to finitness of the limit in Eq. (2), and means that the splash pulse has indeed normal asymptotics, as mentioned already in Sect.1. The same holds also for the real parts of the wave functions in Eqs. (21) and (22), except that the pulse of the real part is bipolar outside of the region of origin at $t=0$, where it is unipolar, and the pulse of the imaginary part is, on the contrary, bipolar. The 50x amplified third curve in Fig. 4 shows the residual tail of the converging wave, which is very weak at $ct=30$ because its peak is already gone by positive times. We shall use Fig. 4 later as a template for graphical illustration for the explanation of the abnormal asymptotics of the primitive in Eq. (3).
Quite analogously to Eqs. (20)-(22), the primitive of the splash pulse $\Psi\left(R,t\right)$ can be decomposed into expanding and converging waves, as can be seen readily from Eq. (3)  [^1] If one integrates $\psi_{+}\left(R,t\right)$ and $\psi_{-}\left(R,t\right)$ over time from $\infty$- ∞ up to any finite value $t$, then the real parts of both $\Psi_{+}\left(R,t\right)$ and $\Psi_{-}\left(R,t\right)$ acquire an additive integration constant $C=-ln(\infty)$ which cancels out in Eq. (23) and therefore does not affect $\Psi\left(R,t\right)$., viz.:
@@
In Fig. 5 the real parts of expanding and converging waves given, respectively, by Eq. (24) and Eq. (25), as well as of $\Psi\left(R,t\right)$ from Eq. (23), are plotted against the radius $R$ at two different instants of time. We no longer deal with the imaginary parts because their asymptotics is not abnormal. This is understandable, because as integrands in Eq. (3), the imaginary parts of $\psi_{+}\left(R,t\right)$ and $\psi_{-}\left(R,t\right)$ decay with $t\rightarrow\infty$ as $t^{-2}$, while the real parts as $t^{-1}$.
The curves are multiplied by $R$ and so become scaled appropriately for their comparison. We observe that the peaks of the real part of $\Psi_{+}(R,t)$ at different time instants are of equal height. This corresponds to finitness of the limit in Eq.(2) and means that the expanding part of the primitive, taken separately, has normal asymptotics. But we also observe, that if the tails of the converging waves are added— according to Eq. (23) with negative sign—the peaks of $R\cdot\Re\left[\Psi\left(R,t\right)\right]$ of the splash pulse grow logarithmically toward $\infty$- ∞. Indeed, the sum of the ordinate of the peak at $R=30$ of $R\cdot\Re\left[\Psi_{+}(R,30)\right]$ and that of $-R\cdot\Re\left[\Psi_{-}(R,30)\right]$ is exactly equal to the ordinate of the peak of $R\cdot\Re\left[\Psi(R,30)\right]$ at the same abscissa $R=30$. The same can be observed for the curves for the time instant $ct=160$.
@@
V Discussion and study of physical feasibility

### v3

@@ Abstract
The behavior of wave signals in the far zone is not only of theoretical interest but also of paramount practical importance in communications and other fields of applications of optical, electromagnetic or acoustic waves. Long time ago T. T. Wu [J. Appl. Phys. 57, 2370 (1985)] introduced models of "electromagnetic missiles" whose decay could be made arbitrarily slower than the usual inverse distance by an appropriate choice of the high frequency portion of the source spectrum. Very recent work by Plachenov and Kiselev [Diff. Eqs. 60, 1634 (2024).] introduced a finite-energy scalar wave solution, different from Wu’s, decaying slower than inversely proportional with the distance. A physical explanation for the unusual asymptotic behavior of the latter will be given in this article. Furthermore, two additional examples of scalar wave pulses characterized by abnormal slow decay in the far zone will be given and their asymptotic behavior will be discussed. A proof of feasibility of acoustic and electromagnetic fields with the abnormal asymptotics will be described.
journal: opticajournalarticletype: Research Article
1 Introduction
It is a textbook truth that solutions of the wave equation decay inversely proportional to distance in the far field or the wave zone (see, e.g., [1, 2]). Long time ago Wu [3] showed that the decay of radiatation from a finite-size antenna can be made arbitrarily slower thanks to stretching the Fresnel region for a part of the source spectrum by a careful choice of the high-frequency wing of the spectrum. However, subsequent experiments did not substantiate claims of unlimited slow decay [4]. A particular study of quasi-missile behavior was undertaken by Shaarawi et al. [5]. They examined in detail the asymptotic behavior of a particular scalar localized wave, referred to as double-exponential pulse (DEX). In the 1980s, closely related to Wu’s "electromagnetic missiles", another subject dealing with uncommon solutions of the wave equation emerged—the so-called localized or non-diffracting or space-time wave packets possessing exotic properties, such as self-healing, superluminality, and invariant propagation without spread or decay, theoretically up to infinity. By now this subject has grown into a large research field of its own, see monograph [6] and reviews [7, 8, 9, 10, 11]. Therefore, the decay in the wave zone as a specific subject in mathematical physics has remained somewhat in the shadows. However, in recent years, several papers on the study of asymptotical behavior have appeared, see [12, 13, 14] and references therein.
The very recent paper by Plachenov and Kiselev [14] (see also [15]) introduced a finite-energy solution, different from Wu’s model [3], that decays slower than inversely proportional to the distance.
They consider the following simple solution
@@ 1 Introduction
to the homogeneous wave equation
$\frac{1}{\rho}\frac{\partial}{\partial\rho}\left(\rho\frac{\partial\psi}{\partial\rho}\right)+\frac{\partial^{2}}{\partial z^{2}}\psi=\frac{\partial^{2}}{\partial(ct)^{2}}\psi\;,$ (2)
where $R=\left(\rho^{2}+z^{2}\right)^{1/2}$ is the distance of the field point from the origin, $c$ is the velocity of light (or sound), and $t_{s}$ is a constant that determines the width of the pulse. Here differently from [14] we have used somewhat changed designations and wrote the wave equation in cylindrical coordinates for the axisymmetric case which is satisfied by the solutions introduced in the following sections.
$\psi\left(R,t\right)$ is a special case of a splash pulse [16]. It is frequently used to model ultrashort pulses and constitutes a spherically propagating pulse, while the spherical surface, where the pulse peaks at $R\approx ct,$ collapses into the origin at negative times $t<0$ and expands for $t>0.$ The real part $\operatorname{Re}\psi\left(R,t\right)$ represents a bipolar pulse and the imaginary part $\operatorname{Im}\psi\left(R,t\right)$ a unipolar one when the pulse moves in the far field where $R\gg ct_{s}$, and vice versa near the origin where $R<ct_{s}$.
@@
The limit of the type of Eq. (3) plays an important role in various fields of mathematical physics and it is known that for finite-energy "well-behaving" solutions of the wave equation the limit $L$ should be finite (incl. $L=0$), see, e.g., [1, 2, 12, 13].
@@
$\begin{split}\Psi\left(R,t\right)=\int_{-\infty}^{t}\psi\left(R,t^{\prime}\right)dt^{\prime}&=\\ =\frac{1}{2cR}\ln\frac{ct+ict_{s}-R}{ct+ict_{s}+R},\end{split}$ (4)
which of course also obeys the wave equation, but in contradistinction to the usual behaviour of fields in the wave zone, the quantity in Eq. (3) is diverging:
@@
They also showed that $\Psi\left(R,t\right)$ has finite total energy, i.e., may not be an unphysical solution. However, both the scalar field energy density and Poynting vector have normal asymptotics in the sense of finite values of the limit of the type in Eq. (3).
The general motivation of our article is to draw attention to the phenomenon of abnormally slow decay, which is quite unexpected, at least in the case of physically feasible waves. Our aim is to find additional examples of wave pulses with unusual asymptotics, and to reveal the causes of their abnormally slow decay. We pay special attention to determining the finiteness of the energy of the found pulses because this property is a crucial one for their physical feasibility In the next section we consider one of such examples—the primitive (antiderivative) of the wave function of a unidirectional pulse. Section 3 is devoted to another example—the fractional splash pulse which already itself, without taking its time integral, decays very slowly in the far zone.
The authors in the short communication [14] neither illustrate graphically the behavior of $\Psi\left(R,t\right)$ nor interpret physically its unusual asymptotics. In Section 4 we try to fill in this gap. In Section 5 we discuss some subtleties of the slow decay of the other pulses studied in Sections 2 and 3 and consider the physical feasibility in optics, electromagnetics, or acoustics of the pulses considered in this study.
2 Hemispherical unidirectional pulse
@@
Similarly to $\psi\left(R,t\right)$ in Eq. (1), the real and imaginary parts of $u\left(R,t\right)$ represent bipolar and unipolar pulses, but in contrast to the spherically symmetric case, in the given case the pulses "ride" on a collapsing-expanding tube-like surface with a hemispherical surface inside.
@@
But how does its primitive behave? The integral over time from -$\infty$ to $t$ was taken first by setting $t_{s}=z_{s}=0$ and then making replacements $z\rightarrow z-iz_{s}$ and $t\rightarrow t+it_{s}$. After some algebra, we found the following expression for the primitive, where $g\left(\rho,t\right)$ is given in Eq. (6).
@@
$U_{1}\left(\rho,z,t\right)$ $\equiv\ln\frac{ct+ict_{s}+h\left(\rho,z\right)}{ct+ict_{s}-h\left(\rho,z\right)},$
$h\left(\rho,z\right)\equiv\sqrt{\rho^{2}+z^{*2}},$
$U_{2}\left(\rho,z,t\right)$ $\equiv\ln\frac{\left[ct^{*}z^{*}+g\left(\rho,t\right)h\left(\rho,z\right)\right]\left[z^{*}-h\left(\rho,z\right)\right]}{\left[ct^{*}z^{*}-g\left(\rho,t\right)h\left(\rho,z\right)\right]\left[z^{*}+h\left(\rho,z\right)\right]},$
where $ct^{*}$ stands for $ict+ct_{s}$ and $z^{*}$ for $z-iz_{s}$. The behavior of $U\left(\rho,z,t\right)$ is depicted in Figs. 1 and 2.
Figure 1: The real and imaginary parts of $U\left(\rho,z,t\right)$ at the instants $ct=0$ and $ct=4$ ($c\equiv 1$). As the plot is axisymmetric, the axis $x$ represents any axis transverse to the propagation axis $z$ and, in distinction from the radial coordinate $\rho,$ also takes negative values. The sign of the imaginary part has been reversed. The pulse width parameters are chosen as follows: $ct_{s}=0.3$ and $z_{s}=0.1$.
Figure 2: "Top view" of the 3D plot of the real part of $U\left(\rho,z,t\right)$ at four instants $t$. See caption of Fig. 1.
We see that the pulse peaks on the surface of a collapsing-expanding tube with a hemisphere inside it, while the radius of both is equal to $|ct|$. Hence, $U\left(\rho,z,t\right)$ behaves similarly to $u\left(\rho,z,t\right)$ from Eq. (6), the most remarkable difference being that the peak of the real part of the latter is abruptly and symmetrically bipolar, except at $t=0$ when it is unipolar contrary to Fig. 1.
@@
Another exception where the limit proves to be finite and equals to $\ln 2$ is when time $t$ runs to -$\infty$ but $z$ to $\infty$. In both these exceptional cases the pulse peak is not involved into the process, as can be seen by looking at Figs. 1 and 2.
We calculated the Poynting vector and energy density of $U\left(\rho,z,t\right)$ according to known formulas [23] (see also Eqs. (1)-(5) of Ref. [21]) for scalar fields, as well as the total energy. The latter is finite. However, both the Poynting vector and the energy density have normal asymptotics in the sense of finite values of the limit of the type in Eq. (3) as they have to be since both consist of derivatives of $U\left(\rho,z,t\right)$ which have normal asymptotics.
Despite the unidirectionality of $U\left(\rho,z,t\right)$—integration as a linear operation preserves the unidirectional spectrum of $u\left(\rho,z,t\right)$— in certain spatial regions the energy flux turned out to be directed opposite to the azis $z$, and at some points the velocity of such an energy backflow even reached almost the value c-c.
@@
Finally, we found that the imaginary part of $U\left(\rho,z,t\right)$ is a "strange" field [24, 25, 26] in the sense that in certain regions the time integral from -$\infty$ to $\infty$ is not zero. This is understandable, because the imaginary part is a unipolar pulse.
3 Fractional splash pulse
It is known that specific members of a general family of splash pulses are derivable as spectral superpositions of the most investigated localized propagation-invariant unipolar pulse—the focus wave mode [7, 8, 6, 16]. A specific fractional splash pulse is given by the following expression
@@ 3 Fractional splash pulse
$\times\frac{1}{\left(a_{2}-i\left(z+ct\right)+\rho^{2}\left(a_{1}+i\left(z-ct\right)\right)^{-1}\right)^{3/4}},$
where $a_{1}$ and $a_{2}$ are positive parameters. The general form of the splash mode pulse is given by the expression in Eq. (14), with the power $3/4$ in the denominator replaced by $\nu+1$, with $\nu>-1$. The fractional pulse above corresponds to $\nu=-1/4$. We note that in the case of DEX pulses [5], which are presented as differences of two splash pulses, the slow decay appears also if $\nu<-1$, but such pulses are not square integrable and the total energy of corresponding fields diverges.
The behavior of $f\left(\rho,z,t\right)$ is depicted in Fig. 3. Both the real and imaginary parts of the pulse possess a collapsing-expanding spherical structure, but a peak appears only in a narrow cone of one hemisphere (of positive $z$ if $ct>0$ and of negative $z$ if $ct<0$). The real part has mirror symmetry with respect to time reversal $t\rightarrow-t$, while the imaginary part additionally reverses its sign. Outside the region of the origin, the pulse tops of the real part and the imaginary part do not coincide: the top of the former is shifted farther from the origin $z=0$ than the point $z=ct$ by a generally small distance about $2a_{1}/3$, and the top of the imaginary part is shifted closer by the distance about $a_{1}/4$. This follows from a detailed inspection of Eq. (14), as well as from plots which will be presented in Section 5. But the most remarkable distinction from the pulses considered in the previous sections is that the fractional splash pulse itself, without the need to take its antiderivative, exhibits abnormal asymptotics as the limits of the type in Eq. (3) show. The limits along different directions turn out to be as follows.
@@
Eq. (15) shows that the peaks of both the real and imaginary parts decay slower than the normal decay $\sim 1/z$ when one moves with the pulse top along the positive direction of the axis $z$ towards infinity. How they approach infinity will be discussed in Section 5. As was mentioned above, the tops of the pulses of the real and imaginary parts of $f\left(\rho,z,t\right)$ are shifted from the point $z=ct$, and therefore the value of $\Delta$ affects differently the limits of the parts. If $ct>0$ and $z>0$, a simple choice is $\Delta=0$ in which case the running point is simultaneously close to the tops of the pulses of both the real and imaginary parts and the exponent in Eq. (15) shows that at this point both $ct\,\Re f\left(\rho,z=ct,t\right)$ and $ct\,\Im f\left(\rho,z=ct,t\right)$ approach infinity in the same way.
@@
We calculated the Poynting vector and energy density of $f\left(\rho,z,t\right)$. Although individually time and space derivatives of the fractional splash pulse exhibit abnormal asymptotic behavior, the energy density and the Poynting vector when inserted into Eq. (15) instead of $f$ do not result in an infinite limit because both quantities are quadratic in time and space derivatives. However, for energy-type quantities, a normal decay is inversely proportional to the square of the distance. This means that in testing the asymptotic behavior, $ct$ must be replaced by $c^{2}t^{2}$ in the expression of the limit. In doing so, we arrive at the conclusion that the energy density and the Poynting vector also decay abnormally slowly. The total energy is finite. $f\left(\rho,z,t\right)$ is square integrable (see Appendix), which is also important from a physical point of view as will be discussed in Section 5.
We studied also other fractional splash pulses which are given by the same expression in Eq. (14) but with other values of the power instead of $3/4$ in the denominator. If the power is $1/2$, the pulse has the same properties as $f\left(\rho,z,t\right)$, but unfortunately its total energy is infinite. But if the value of the power is between $1/2$ and $1$, i.e., if $\frac{1}{2}<\nu+1<1$, the pulse has slower than normal decay and at the same time its wave function is square-integrable and has finite total energy (see Appendix).
4 Causes of uncommon asymptotical behavior at infinity
@@
$\rho(\textbf{r},t)=\delta(\textbf{r})\hskip 2.0pt\frac{1}{2c(t+it_{s})}~,$
@@
Figure 4: Radial dependence of the imaginary parts of the expanding and converging waves defined in Eq. (21) and Eq. (22) at instants $ct=30$, $ct=170$, and ct=100ct=-100. All curves have been multiplied by R-R and the third curve additionally by 50. The pulsewidth parameter $ct_{s}=2$.
In Fig. 4 the imaginary parts of the expanding and converging waves given by Eq. (21) and Eq. (22), respectively, are plotted against the radius $R$ at three different instants of time. The curves are plotted with reverse sign for better readability and are scaled—multiplied by RR—in order to eliminate the factor $1/R$. Thanks to the scaling we observe that the peaks of the pulses at different time instants are of equal height. This correponds to finitness of the limit in Eq. (3), and means that the splash pulse has indeed normal asymptotics, as mentioned already in Section 1. The same holds also for the real parts of the wave functions in Eqs.  (21) and (22), except that the pulse of the real part is bipolar outside of the region of origin at $t=0$, where it is unipolar, and the pulse of the imaginary part is, on the contrary, bipolar. The 50x amplified third curve in Fig. 4 shows the residual tail of the converging wave, which is very weak at $ct=30$ because its peak is already gone by positive times. We shall use Fig. 4 later as a template for graphical illustration for the explanation of the abnormal asymptotics of the primitive in Eq. (4).
Quite analogously to Eqs. (20)-(22), the primitive of the splash pulse $\Psi\left(R,t\right)$ can be decomposed into expanding and converging waves, as can be seen readily from Eq. (4)  [^1] If one integrates $\psi_{+}\left(R,t\right)$ and $\psi_{-}\left(R,t\right)$ over time from -$\infty$ up to any finite value $t$, then the real parts of both $\Psi_{+}\left(R,t\right)$ and $\Psi_{-}\left(R,t\right)$ acquire an additive integration constant $C=-ln(\infty)$ which cancels out in Eq. (23) and therefore does not affect $\Psi\left(R,t\right)$., viz.:
@@
In Fig. 5 the real parts of expanding and converging waves given, respectively, by Eq. (24) and Eq. (25), as well as of $\Psi\left(R,t\right)$ from Eq. (23), are plotted against the radius $R$ at two different instants of time. We no longer deal with the imaginary parts because their asymptotics are not abnormal. This is understandable because, as integrands in Eq. (4), the imaginary parts of $\psi_{+}\left(R,t\right)$ and $\psi_{-}\left(R,t\right)$ decay with $t\rightarrow\infty$ as $t^{-2}$, while the real parts as $t^{-1}$.
Figure 5: Radial dependencies of the real parts of the waves defined in Eq.s (24), (25), and (23) at instants $ct=30$ and $ct=160$. All curves have been multiplied by $R$. The pulse width parameter is $ct_{s}=2$. For comparison, a logarithmically diverging dependence is shown by the dotted curve, whereas the factor $1/2$ has been inserted in accordance with the presence of the same factor in Eqs. (24) and (25).
The curves are multiplied by $R$ and thus are appropriately scaled for comparison. We observe that the peaks of the real part of $\Psi_{+}(R,t)$ at different time instants are of equal height. This corresponds to the finitness of the limit in Eq.(3) and means that the expanding part of the primitive, taken separately, has normal asymptotics. But we also observe that if the tails of the converging waves are added— according to Eq. (23) with negative sign—the peaks of $R\cdot\Re\left[\Psi\left(R,t\right)\right]$ of the splash pulse grow logarithmically toward -$\infty$. Indeed, the sum of the ordinate of the peak at $R=30$ of $R\cdot\Re\left[\Psi_{+}(R,30)\right]$ and that of $R\cdot\Re\left[\Psi_{-}(R,30)\right]$ is exactly equal to the ordinate of the peak of $R\cdot\Re\left[\Psi(R,30)\right]$ at the same abscissa $R=30$. The same can be observed for the curves for the time instant $ct=160$.
@@
5 Discussion and study of physical feasibility